OpenStudy (anonymous):

What is the number of real solutions? 6x^2 – 7x – 1 = 0 two solutions cannot be determined one solution no real solutions

myininaya (myininaya):

Find the discriminant. \[\text{ If you have } ax^2+bx+c=0 \text{ , then the discriminant is equal to } D= b^2-4ac .\] If D=0, then one real solution (with multiplicity 2) If D=negative real number, then no real solutions. If D=positive real number, then 2 real solutions.

OpenStudy (anonymous):

I got a dismals and i think its A.

OpenStudy (anonymous):

but i wanna make sure i got it right,im struggling with math

myininaya (myininaya):

I don't know what dismals mean but yes D would be a positive real number which implies 2 real solutions. You have a=6,b=-7,c=-1 You put these in the formula I gave above for the discriminant.

OpenStudy (anonymous):

Tip: when "a" and "c" are of opposite signs, you will always have real solutions.